OpenStudy (anonymous):

Determine whether or not the improper integral converges:

OpenStudy (anonymous):

\[\int\limits_{4}^{\infty}1/x \sqrt{x} \] dx

OpenStudy (anonymous):

Moya, can you integrate that from 4 to some variable, say "t"?

OpenStudy (kropot72):

Is this it? \[\int\limits_{4}^{\infty}\frac{1}{x \sqrt{x}}\]

OpenStudy (anonymous):

kropot2 made a good point. Please clarify the integrand.

OpenStudy (anonymous):

yes thats it kropot. @ gpfreitas isnt it the infity thtat is replaced with t instead of the 4?

OpenStudy (anonymous):

@Moya. Yes, you should first replace the infinity by t. That is why I said "from 4 to t".

OpenStudy (anonymous):

Ok, what is the indefinite integral of \[\frac{1}{x \sqrt{x}}\] ?

OpenStudy (anonymous):

Hint: write this integrand as \(x^k\) for some fraction \(k\).

OpenStudy (anonymous):

should i rewrite the funtion so it becomes a product that is lnx*x^-1/2??

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

Simpler than that. \(1/x\) is \(x^{-1}\). \(1/\sqrt{x}\) is \(x^{-1/2}\). The integrand is the product of those two expressions, which is...

OpenStudy (anonymous):

Anyway, you will get something that is of the form \(x^{k}\) where \(k\) will be some fraction. I am guessing you know how to integrate something of that form. Integrate it from 4 to \(t\). You will get a formula that depends only on \(t\). Take the limit of that formula when \(t \to \infty\). That is the answer. Do you see what you have to do?

OpenStudy (anonymous):

x^1/2

OpenStudy (anonymous):

Hmmmm... are you sure? Check your calculation again. Make sure you got the signs right.

OpenStudy (anonymous):

oh x^-3/2

OpenStudy (anonymous):

Yep. Next step?

OpenStudy (anonymous):

(Quick aside; try to get into the habit of writing x^(-3/2), with parentheses, or use the math notation \(x^{-3/2}\). What you wrote could be interpreted as \[\frac{x^{-3}}{2}\] which is not what you meant.)

OpenStudy (anonymous):

thanks i think i have the idea now, ill make x tend to t and substitute in the t and 4 then find the limit of that answer right?

OpenStudy (anonymous):

I'm not sure I understand you. To solve it, you don't really make \(x\) tend to \(t\)... You integrate \(x^{-3/2}\) from 4 to some variable t. From that you will get an expression that depends on \(t\), but not on \(x\). Take the limit of that when \(t \to \infty\). That is all. Sorry, but I need to go now! Good luck!